40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1526     Entry time: Tue Sep 23 18:40:08 2014
Author: Evan 
Type: DailyProgress 
Category: EOM 
Subject: No more 2 kHz hump in beat 

Background

Yesterday I think I narrowed down the source of the 2 kHz frequency noise hump: it is voltage noise from the TTFSS being injected into the broadband EOM.

With the north cavity unlocked (and the TTFSS set to "test"), I monitored the (undemodulated) RAM using the auxiliary 1811 and the HP4395A. There were clear, broad 600 Hz humps on either side of the 14.75 MHz carrier. It disappeared when I unplugged the drive to the broadband EOM.

Then I looked at various test points on the TTFSS HV board with the SR785. On the COM → EOM path, the TF shaping takes the COM noise and produces (what I think is) the same 600 Hz bump, which is then sent to the EOM. In the beat, the bump appears at 2 kHz because of the north TTFSS boost; with the boost off, it reverts to 600 Hz.

This is the case on both TTFSS boards, but it only leaked into the beat on the north cavity. So I suspected it was an issue with how the EOMs are aligned on the north path. On north, the BB EOM was immediately followed by the resonant PDH EOM; on south, between the BB EOM and PDH EOM there is a PMC, an FI, and some other optics.

Today's work

I moved the resonant EOM so that it follows the EOAM. After the post-EOAM PBS, I did the following:

  • I set down a HWP, and then used a temporary PBS to ensure s-polarization of the beam.
  • A few inches after the first HWP, I set down a second HWP and used a temporary PBS to ensure p-polarization of the beam.
  • Between the HWPs, I placed the resonant EOM, screwed it down, and then aligned the beam through it.

Then I redid the mode-matching into the north cavity and measured the beat. I kept it locked for about 90 minutes and didn't see the 2 kHz hump appear, so I'm guessing this solved the issue.

To do

  • Minimize RAM on north cavity
  • RIN data is stale and needs to be retaken
  • Need to fix a nominal operating power for beat PD (I pick 7 dBm, because we're using a ZRPD-1 phase detector)
  • Marconi noise data is stale and needs to be retaken
  • PLL readout data is stale and needs to be retaken
  • Seismic data is stale and needs to be retaken
Attachment 1: noiseBudget.pdf  1.086 MB  Uploaded Tue Sep 23 19:41:11 2014  | Hide | Hide all
noiseBudget.pdf
ELOG V3.1.3-